OpenStudy (anonymus11):

Find a quadratic equation with roots -1+4i and -1-4i

OpenStudy (anonymus11):

Algebra 2 please help

OpenStudy (imstuck):

I got the quadratic equation to be \[x ^{2}+2x-15\]

OpenStudy (imstuck):

Let me check it, then I'll explain it, ok?

OpenStudy (anonymus11):

Ok thank you!!

OpenStudy (imstuck):

Ok, I believe that that is correct; it's been a long time since I taught this!!! The roots that we are considering are x = -1+4i and x = -1-4i. But in order to go backwards from that to the original quadratic, we have to have (x +/- something)(x +/- something) to get an x^2 term, cuz quadratics ARE x^2 terms. So if x = -1+4i, then the root would be (x + 1 + 4i). And the other root would be (x + 1 - 4i). So here's what we have: \[(x+1+4i)(x+1-4i)\]

OpenStudy (imstuck):

Now we FOIL that out, starting with the first x in the first set of parenthesis: \[x ^{2}+x-4ix\]

OpenStudy (imstuck):

Now let's move to the 1 in the first set of parenthesis: \[x+1-4i\]

OpenStudy (volleyballlover55):

@IMStuck hi! when your done please come back to my question :)

OpenStudy (imstuck):

Now let's move to the 4i in the first set of parenthesis: \[4ix+4i-16i ^{2}\]

OpenStudy (imstuck):

Now let's put it all together: \[x ^{2}+x-4ix+x+1-4i+4ix+4i-16i ^{2}\]

OpenStudy (imstuck):

Combining like terms and canceling out the ones that eliminate each other, we are left with: \[x ^{2}+2x-16i ^{2}+1\]

OpenStudy (imstuck):

The thing you need to remember here is that i^2 = -1: \[x ^{2}+2x-16(-1)+1=x ^{2}+2x+16+1=x ^{2}+2x+17\]

OpenStudy (imstuck):

I did my negatives incorrectly when I gave you my initial equation. A 2x -(-16) is 2x + 16. My bad.

OpenStudy (imstuck):

So the equation, as far as I can possibly remember, is \[x ^{2}+2x+17\]

OpenStudy (imstuck):

I am going to check that now using the quadratic formula...will return in just a moment!

OpenStudy (imstuck):

Yes, that is correct. Good luck with those! They are no different than FOILing out regular roots.

OpenStudy (zzr0ck3r):

Another way of looking at it. \(x=-1\pm4i\implies x+1=\pm4i\implies (x+1)^2=(4i)^2\) Which gives \(x^2+2x+17=0\).